# Entropy Functions on Two-Dimensional Faces of Polymatroidal Region of Degree Four: Part II: Information Theoretic Constraints Breed New Combinatorial Structures

### v1

@@ III-A Entropy functions on faces with extreme rays containing $U_{2,4}$ and one rank-1 matroid
which implies $p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4})=0$ contradicting $p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4})>0$. Similarly, we can show that by (III.1.101), $p_{X_{1},X_{2},X_{3},X_{4}}(1,1,0,x_{4})>0$ will also lead a contradiction.
@@
Remark: In this theorem, we give an inner bound on the face $(U_{2,4},U^{4}_{1,1})$. Entropy functions on this face corresponds to a pair orthogonal squares, one is Latin and the other is muti-symbol Latin. A square is called a multi-symbol Lain square of order $v$ with symbol set size $v^{\prime}\geq v$ if it is a $v\times v$ square with a set of of symbols with size $v^{\prime}$ and each cell contains one or more symbols, and each symbol appears in each row and each column exactly once. Such kind of pair of squares can be obtainded by splitting the symbols of one square of a pair of orthogonal Latin squares for $v\neq 2,6$. We proved that such pair does not exist for $v=2$. For $v=6$, we gave a pair with $v^{\prime}=12$. We conjecture that this inner bound is tight.
III-B Entropy functions on faces with extreme rays both containing rank $2$ integer polymatroids
In this subsection, we characterize entropy functions on three $2$-dimensional $(U_{2,4},\mathcal{W}^{12}_{2})$ faces $(U_{2,4},U^{123}_{2,3})$, and $(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$ of $\Gamma_{4}$ with extreme rays both containing rank $2$ interger polymatroids. Some other faces in this family $(U_{2,3}^{123},U_{2,3}^{124})$ and $(\mathcal{W}_{2}^{12},U_{2,3}^{134})$ have already been characterized in Part I of this serize of two papers, while $(\hat{U}_{2,5}^{1},U^{234}_{2,3})$ will be characterized in Subsection III-C as Latin square decomposition will be used.
@@ III-B Entropy functions on faces with extreme rays both containing rank $2$ integer polymatroids
Remark: This theorem studies the relations of two Latin squares (of first order) of the same size from the perspective of entropy functions, which generalize both two orthogonal Latin squares and two identical Latin squares. For any pair of two $v\times v$ Latin squares $S_{1}$ and $S_{2}$, it correspond to a $v^{2}\times 4$ $\mathbf{T}$ array with each row $(i,j,s_{1},s_{2})$, where $i,j\in\mathbb{I}_{v}$ and $s_{1}$ and $s_{2}$ are the symbols in $S_{1}(i,j)$ and $S_{2}(i,j)$, respectively. On one hand, when $S_{1}$ and $S_{2}$ are orthogonal, entropy function $\mathbf{h}$ of random vector distributed on the rows of $\mathbf{T}$ is in the extreme ray containing $U_{2,4}$; on the other hand, when they are identical, $\mathbf{h}$ is in the extreme ray containing $\mathcal{W}^{12}_{2}$. When they are neither orthogonal, or identical, $\mathbf{h}$ locates in the face but not in the extreme ray.
@@
and the choose of $(x_{2},x_{3})\in\mathcal{X}_{2}\times\mathcal{X}_{3}$ can be arbitrary for a fixed $(x_{1},x_{4})\in\mathcal{X}_{1}\times\mathcal{X}_{4}$, replacing $p(x_{1},x^{\prime}_{2},x^{\prime}_{3},x_{4})$ by $\frac{1}{v^{2}}$ in (III.2.29), we obtain
@@
  • For any positive integer $v\neq 2,6$, let $\bm{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{t})$ be a partition of $k$, and $\{A_{i},i=1,\dots,t\}$ be a partition of $\mathbb{I}_{v}$ with $|A(i)|=\alpha_{i}$. Let $\mathbf{T}$ be a $\mathrm{VOA}(U_{2,4},v)$, and $\mathbf{T^{\prime}}$ be a $v^{2}\times 4$ array such that $\mathbf{T}^{\prime}(i)=\mathbf{T}(i)$ for $i=1,2,3$ and each entry in $\mathbf{T}^{\prime}(4)$ be $j$ if the corresponding entry in $\mathbf{T}(4)$ is in $A_{j}$. Then let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T^{\prime}}.$ It can be checked that $(X_{i},i\in N_{4})$ characterizes $(a,b)$.
@@
For $v=2$ and $6$, as there is no such $\mathrm{VOA}(U_{2,4},2)$ or $\mathrm{VOA}(U_{2,4},6)$ [6, Proposition 2], the above construction is invalid, and so we have to discuss them separately.
@@
Remark: A frequency square induced by partition $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{t})$ of integer $v$ is a $v\times v$ square $S$ with symbols $k\in\mathbb{I}_{t}$ each appearing in each row and each column of $S$ $\alpha_{k}$ times. Note that when $\bm{\alpha}$ is the all-$1$ partition of $v$, $S$ reduces to a Latin square. A Latin square $S_{1}$ of order $v$ (and of first order) and a frequency square $S_{2}$ induced by partition $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{t})$ of integer $v$ are called orthogonal if each pair $(s_{1},s_{2})\in\mathbb{I}_{v}\times\mathbb{I}_{t}$ appears $\alpha_{i}$, where $s_{2}$ is the symbol appear $\alpha_{i}$ times in each row and column of $S_{2}$. By characterizing entropy functions on the face $(U_{2,4},U^{123}_{2,3})$, this theorem also studies arrays corresponding to orthogonal two squares with one Latin square $S_{1}$ and one frequency square $S_{2}$. When $\bm{\alpha}$ is the all-$1$ partition of $v$, it reduces to orthogonal of two Latin squares and then corresponding entropy functions are on the extreme ray containing $U_{2,4}$. On the other hand, when $\bm{\alpha}$ is the trivial partition of single $v$, it reduced to orthogonal squares of one Latin square of first order and the other of zeroth order, and entropy functions are in the extreme ray containing $U^{123}_{2,3}$. It is interesting that although $\mathrm{VOA}(U_{2,4},6)$ does not exist, arrays “between” $\mathrm{VOA}(U^{123}_{2,3},6)$ and $\mathrm{VOA}(U_{2,4},6)$ exist.
@@
For $F=(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log v$ for some positive $v$ and $a=\frac{1}{v}\sum_{i=0}^{v-1}H(\bm{\alpha_{i}})$, where $\bm{\alpha_{i}}\in\mathcal{P}(v),i\in\mathbb{I}_{v}$.
@@
which implies $X_{2}$ is a function of $X_{1}$. Then for each $j\in\mathcal{X}_{1}$, there exists a unique $i\in\mathcal{X}_{2}=\mathbb{I}_{v}$ such that $(j,i)$ forms a row in $\mathbf{T}(1,2)$. Let $\beta_{i,j}$ denote the times of the row $(j,i)$ that occurs in $\mathbf{T}(1,2)$. Note that each $i\in\mathbb{I}_{v}$ occurs in $\mathbf{T}(2)$ exactly $v$ times by the definition of a $\mathrm{VOA}(U_{2,3},k)$. So for each $i\in\mathbb{I}_{v}$, $\bm{\beta_{i}}=(\beta_{i,j}>0,j\in\mathcal{X}_{1})$ forms a partition of $v$. We assume that there exist $t_{i}$ different $j$ such that $\beta_{i,j}>0$ for $i\in\mathbb{I}_{v}$. Then, $\bm{\beta_{i}}$ can be written as $\bm{\alpha_{i}}=(\alpha_{i,1},\dots,\alpha_{i,t_{i}})$. Then
@@
$\frac{\alpha_{v-1,0}}{v^{2}},\frac{\alpha_{v-1,2}}{v^{2}},\dots,\frac{\alpha_{v-1,t_{v}-1}}{v^{2}})$
@@
Now we prove the “if” part. For any $\mathbf{h}=(a,b)$ satisfying $a+b=\log v$ and $a=\frac{1}{v}\sum_{i=1}^{v}H(\bm{\alpha_{i}})$, where $\bm{\alpha_{i}}=(\alpha_{i,1},\alpha_{i,2},\dots,\alpha_{i,t_{i}})\in\mathcal{P}(v),i=1,2,...,v$, let $\mathbf{T}$ be a $v^{2}\times 4$ array such that $\mathbf{T}(2,3,4)$ is a $\mathrm{VOA}(U_{2,3},v)$. In the rows where $i$ occurs in the $\mathbf{T}(2)$, let the entry $(\sum_{m=0}^{i-1}t_{m})+j$ occur $\alpha_{ij}$ times in $\mathbf{T}(1)$ for $i\in\mathbb{I}_{v}$ and $j\in\mathbb{I}_{t_{i}}$ where we set $t_{-1}=0$. Then let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T}.$ It can be checked that $(X_{i},i\in N_{4})$ characterizes $(a,b)$. The proof is accomplished. ∎
Remark: Similar to the first two faces in this Subsection, the characterization of $(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$ can also be considered as two squares of order $v$, where $S_{2}$ is a Latin square and $S_{1}$ is a square whose symbols can be obtained from splitting symbols in $S_{2}$, that is, each symbol $i$ splits into $t_{i}$ symbols with each occurs $\alpha_{i,j}$ times according to the partition $\bm{\alpha}_{i}$.
@@ III-C Entropy functions on faces involving Latin square decompositions
$a=\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|B_{i}|\quad\text{ and }\quad b=\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|A_{i}|,$
@@
$H(X_{1},X_{2})$ $=H(\underbrace{\dfrac{|B_{0}|}{v^{2}},\dots,\dfrac{|B_{0}|}{v^{2}}}_{|A_{0}|},\underbrace{\dfrac{|B_{1}|}{v^{2}},\dots,\dfrac{|B_{1}|}{v^{2}}}_{|A_{1}|},\dots,\underbrace{\dfrac{|B_{v-1}|}{v^{2}},\dots,\dfrac{|B_{v-1}|}{v^{2}}}_{|A_{v-1}|})$ (III.3.20)
@@
Remark: Theorem III-C.1 establish a correspondence between the the $2$-dim face characterization problem and uniform decomposition problem of a $\mathrm{VOA}(U_{2,3},v)$. When $k$ is prime, $\mathrm{VOA}(U_{2,3},v)$ can be decomposed into $k$ uniform subarrays where either $|A_{i}|=1$ and $|B_{i}|=v$ for $i\in\mathbb{I}_{v}$, or $|A_{i}|=v$ and $|B_{i}|=1$ for $i\in\mathbb{I}_{v}$. These correspond to the polymatroids $(0,\log v)$ and $(\log v,0)$, respectively. While for a composite $v$, the uniform decomposition of a $\mathrm{VOA}(U_{2,3},v)$ with be more complicated. In Example 4, the uniform decomposition corresponds to the entropy function $(0.5,1.5)$ on the face $F$.
@@
$a=\frac{1}{2}H(\dfrac{|A_{i}|^{2}}{v^{2}}:i\in\mathbb{I}_{t}).$ (III.3.52)
@@
Remark: Theorem III-C.2 establishes a correspondce beween the face $(\hat{U}_{2,5}^{1},U^{234}_{2,3})$ characterization and suborder decomposition problem. It is obvious that $\mathrm{VOA}(U_{2,3},v)$ is inherently a suborder $\mathrm{VOA}$ of itself. On the other hand, any $\mathrm{VOA}(U_{2,3},v)$ can be decomposed into $v^{2}$ suborder VOA $\mathrm{VOA}(U_{2,3},1)$. These two cases correspond to the polymatroids $(0,\log v)$ and $(\log v,0)$, respectively. However, listing all the $\mathrm{VOA}$ decompositions of a $\mathrm{VOA}(U_{2,3},v)$ can be challenging.
@@
  • for each $i=2,3,4$, each row in $\mathbf{T}(1,i)$ occurs exactly onces,
@@
$p(x_{1},x_{2},x_{3},x_{4})$ $=p(x_{1},x_{2},x_{3})$ (III.3.53)
$=p(x_{1},x_{2},x_{4})$ (III.3.54)
$=p(x_{1},x_{3},x_{4})$ (III.3.55)
$=p(x_{2},x_{3},x_{4})$ (III.3.56)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.3.57)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.3.58)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.3.59)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1},x_{2},x_{3},x_{4}),$ (III.3.60)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1},x_{2},x_{3},x_{4}),$ (III.3.61)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1},x_{2},x_{3},x_{4}).$ (III.3.62)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1},x_{2},x_{3},x_{4}),$ (III.3.63)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1},x_{2},x_{3},x_{4}),$ (III.3.64)
$p(x_{1},x_{3},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1},x_{2},x_{3},x_{4}).$ (III.3.65)
By (III.3.53), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (III.3.60), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{2}).$ (III.3.66)
Together with (III.3.54), we obtain
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2}).$ (III.3.67)
@@
By (III.3.56), $p(x_{1},x_{2},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$, which implies that $(X_{i},i\in N_{4})$ must be uniformly distributed on the rows of a $v^{2}\times 4$ array $\mathbf{T}$ with $\mathbf{T}(2,3,4)$ a $\mathrm{VOA}(U_{2,3},v)$. Note that $p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})$ by (III.3.67), each row of $\mathbf{T}(1,2)$ occurs exactly once in $\mathbf{T}(1,2)$. Similarly, by (III.3.68), each row of $\mathbf{T}(A)$ occurs exactly once in $\mathbf{T}(A)$ for $A=\{1,3\}$ and $M=\{1,4\}$. Hence, $\mathbf{T}$ is a $\{1\}$-partial $\mathrm{VOA}(U_{2,4})$. Recall that $(X_{i},i\in N_{4})$ is uniformly distributed on the rows of $\mathbf{T}$, the probability of each row $(x_{1},x_{2},x_{3},x_{4})$ of $\mathbf{T}$ is
$p(x_{1},x_{2},x_{3},x_{4})=\frac{1}{v^{2}}.$ (III.3.69)

### v2

@@ III-A Entropy functions on faces with extreme rays containing $U_{2,4}$ and one rank-1 matroid
which implies $p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4})=0$ contradicting $p_{X_{1}X_{2}X_{3}X_{4}}(1,0,1,x_{4})>0$. Similarly, we can show that by (III.1.101), $p_{X_{1}X_{2}X_{3}X_{4}}(1,1,0,x_{4})>0$ will also lead a contradiction.
@@
Remark In this theorem, we give an inner bound on the face $(U_{2,4},U^{4}_{1,1})$. Entropy functions on this face corresponds to a pair orthogonal squares, one is Latin and the other is multi-symbol Latin. A square is called a multi-symbol Lain square of order $v$ with symbol set size $v^{\prime}\geq v$ if it is a $v\times v$ square with a set of symbols with size $v^{\prime}$ and each cell contains one or more symbols, and each symbol appears in each row and each column exactly once. Such kind of pair of squares can be obtained by splitting the symbols of one square of a pair of orthogonal Latin squares for $v\neq 2,6$. We proved that such a pair does not exist for $v=2$. For $v=6$, we gave a pair with $v^{\prime}=12$. We conjecture that this inner bound is tight.
III-B Entropy functions on faces with both extreme rays containing rank $2$ integer polymatroids
In this subsection, we characterize entropy functions on three $2$-dimensional faces $(U_{2,4},\mathcal{W}^{12}_{2})$, $(U_{2,4},U^{123}_{2,3})$, and $(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$ of $\Gamma_{4}$ with extreme rays both containing rank $2$ integer polymatroids. Some other faces in this family $(U_{2,3}^{123},U_{2,3}^{124})$ and $(\mathcal{W}_{2}^{12},U_{2,3}^{134})$ have already been characterized in Part I of this series of two papers, while $(\hat{U}_{2,5}^{1},U^{234}_{2,3})$ will be characterized in Subsection III-C as Latin square decomposition will be used.
@@ III-B Entropy functions on faces with both extreme rays containing rank $2$ integer polymatroids
Remark This theorem studies the relations of two Latin squares (of first order) of the same size from the perspective of entropy functions, which generalize both orthogonal Latin squares and two identical Latin squares. For any pair of two $v\times v$ Latin squares $S_{1}$ and $S_{2}$, it correspond to a $v^{2}\times 4$ $\mathbf{T}$ array with each row $(i,j,s_{1},s_{2})$, where $i,j\in\mathbb{I}_{v}$ and $s_{1}$ and $s_{2}$ are the symbols in $S_{1}(i,j)$ and $S_{2}(i,j)$, respectively. On one hand, when $S_{1}$ and $S_{2}$ are orthogonal, entropy function $\mathbf{h}$ of a random vector distributed on the rows of $\mathbf{T}$ is in the extreme ray containing $U_{2,4}$; on the other hand, when they are identical, $\mathbf{h}$ is in the extreme ray containing $\mathcal{W}^{12}_{2}$. When they are neither orthogonal nor identical, $\mathbf{h}$ locates in the face, but not in the extreme ray.
@@
and the choice of $(x_{2},x_{3})\in\mathcal{X}_{2}\times\mathcal{X}_{3}$ can be arbitrary for a fixed $(x_{1},x_{4})\in\mathcal{X}_{1}\times\mathcal{X}_{4}$, replacing $p(x_{1},x^{\prime}_{2},x^{\prime}_{3},x_{4})$ by $\frac{1}{v^{2}}$ in (III.2.29), we obtain
@@
  • For any positive integer $v\neq 2,6$, let $\bm{\alpha}=(\alpha_{1},\alpha_{2},\cdots,\alpha_{t})$ be a partition of $v$, and $\{A_{i},i=1,\dots,t\}$ be a partition of $\mathbb{I}_{v}$ with $|A(i)|=\alpha_{i}$. Let $\mathbf{T}$ be a $\mathrm{VOA}(U_{2,4},v)$, and $\mathbf{T^{\prime}}$ be a $v^{2}\times 4$ array such that $\mathbf{T}^{\prime}(i)=\mathbf{T}(i)$ for $i=1,2,3$ and each entry in $\mathbf{T}^{\prime}(4)$ be $j$ if the corresponding entry in $\mathbf{T}(4)$ is in $A_{j}$. Then let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T^{\prime}}.$ It can be checked that $(X_{i},i\in N_{4})$ characterizes $(a,b)$.
@@
For $v=2$ and $6$, as there is no such $\mathrm{VOA}(U_{2,4},2)$ or $\mathrm{VOA}(U_{2,4},6)$ [7, Proposition 2], the above construction is invalid, and so we have to discuss them separately.
@@
Remark A frequency square induced by partition $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{t})$ of integer $v$ is a $v\times v$ square $S$ with symbols $k\in\mathbb{I}_{t}$, each appearing in each row and each column of $S$ $\alpha_{k}$ times. Note that when $\bm{\alpha}$ is the all-$1$ partition of $v$, $S$ reduces to a Latin square. A Latin square $S_{1}$ of order $v$ (and of the first kind) and a frequency square $S_{2}$ induced by partition $\bm{\alpha}=(\alpha_{1},\ldots,\alpha_{t})$ of integer $v$ are called orthogonal if each pair $(s_{1},s_{2})\in\mathbb{I}_{v}\times\mathbb{I}_{t}$ appears $\alpha_{i}$, where $s_{2}$ is the symbol that appears $\alpha_{i}$ times in each row and column of $S_{2}$. By characterizing entropy functions on the face $(U_{2,4},U^{123}_{2,3})$, this theorem also studies arrays corresponding to orthogonal two squares with one Latin square $S_{1}$ and one frequency square $S_{2}$. When $\bm{\alpha}$ is the all-$1$ partition of $v$, it reduces to orthogonal of two Latin squares and then corresponding entropy functions are on the extreme ray containing $U_{2,4}$. On the other hand, when $\bm{\alpha}$ is the trivial partition of a single $v$, it reduces to orthogonal squares of one Latin square of the first kind and the other of the zeroth kind, and entropy functions are in the extreme ray containing $U^{123}_{2,3}$. It is interesting that although $\mathrm{VOA}(U_{2,4},6)$ does not exist, arrays “between” $\mathrm{VOA}(U^{123}_{2,3},6)$ and $\mathrm{VOA}(U_{2,4},6)$ exist.
@@
For $F=(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$, $\mathbf{h}=(a,b)\in F$ is entropic if and only if $a+b=\log v$ for some positive $v$ and $a=\frac{1}{v}\sum_{i=0}^{v-1}H(\bm{\alpha_{i}})$, where $\bm{\alpha_{i}}\in\mathcal{P}(v),i\in\mathbb{I}_{v}$. Here, $\mathcal{P}(v)$ denotes the family of all partitions of $v$.
@@
which implies $X_{2}$ is a function of $X_{1}$. Then for each $j\in\mathcal{X}_{1}$, there exists a unique $i\in\mathcal{X}_{2}=\mathbb{I}_{v}$ such that $(j,i)$ forms a row in $\mathbf{T}(1,2)$. Let $\beta_{i,j}$ denote the times of the row $(j,i)$ that occurs in $\mathbf{T}(1,2)$. Note that each $i\in\mathbb{I}_{v}$ occurs in $\mathbf{T}(2)$ exactly $v$ times by the definition of a $\mathrm{VOA}(U_{2,3},k)$. So for each $i\in\mathbb{I}_{v}$, $\bm{\beta_{i}}=(\beta_{i,j}>0,j\in\mathcal{X}_{1})$ forms a partition of $v$. We assume that there exist $t_{i}$ different $j$ such that $\beta_{i,j}>0$ for $i\in\mathbb{I}_{v}$. Then, $\bm{\beta_{i}}$ can be written as $\bm{\alpha_{i}}=(\alpha_{i,0},\dots,\alpha_{i,t_{i}-1})$. Then
@@
$\frac{\alpha_{v-1,0}}{v^{2}},\frac{\alpha_{v-1,1}}{v^{2}},\dots,\frac{\alpha_{v-1,t_{v}-1}}{v^{2}})$
@@
Now we prove the “if” part. For any $\mathbf{h}=(a,b)$ satisfying $a+b=\log v$ and $a=\frac{1}{v}\sum_{i=0}^{v-1}H(\bm{\alpha_{i}})$, where $\bm{\alpha_{i}}=(\alpha_{i,0},\dots,\alpha_{i,t_{i}-1})\in\mathcal{P}(v),i\in\mathbb{I}_{v}$, let $\mathbf{T}$ be a $v^{2}\times 4$ array such that $\mathbf{T}(2,3,4)$ is a $\mathrm{VOA}(U_{2,3},v)$. In the rows where $i$ occurs in the $\mathbf{T}(2)$, let the entry $(\sum_{m=0}^{i-1}t_{m})+j$ occur $\alpha_{i,j}$ times in $\mathbf{T}(1)$ for $i\in\mathbb{I}_{v}$ and $j\in\mathbb{I}_{t_{i}}$ where we set $t_{-1}=0$. Then let $(X_{i},i\in N_{4})$ be uniformly distributed on the rows of $\mathbf{T}.$ It can be checked that $(X_{i},i\in N_{4})$ characterizes $(a,b)$. The proof is accomplished. ∎
Remark Similar to the first two faces in this Subsection, the characterization of $(\hat{U}_{2,5}^{1},\mathcal{W}_{2}^{12})$ can also be considered as two squares of order $v$, where $S_{2}$ is a Latin square and $S_{1}$ is a square whose symbols can be obtained from splitting symbols in $S_{2}$, that is, each symbol $i$ splits into $t_{i}$ symbols with each occurring $\alpha_{i,j}$ times according to the partition $\bm{\alpha}_{i}$.
@@ III-C Entropy functions on faces involving Latin square decompositions
$a=\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|B_{i}|\text{ and }b=\log v-\frac{1}{v}\sum_{i=0}^{v-1}\log|A_{i}|,$
@@
$H(X_{1},X_{2})$ $=H(\underbrace{\dfrac{|B_{0}|}{v^{2}},\dots,\dfrac{|B_{0}|}{v^{2}}}_{|A_{0}|},\dots,\underbrace{\dfrac{|B_{v-1}|}{v^{2}},\dots,\dfrac{|B_{v-1}|}{v^{2}}}_{|A_{v-1}|})$ (III.3.20)
@@
Remark Theorem III-C.1 establishes a correspondence between the $2$-dim face characterization problem and the uniform decomposition problem of a $\mathrm{VOA}(U_{2,3},v)$. When $v$ is prime, $\mathrm{VOA}(U_{2,3},v)$ can be decomposed into $v$ uniform subarrays where either $|A_{i}|=1$ and $|B_{i}|=v$ for $i\in\mathbb{I}_{v}$, or $|A_{i}|=v$ and $|B_{i}|=1$ for $i\in\mathbb{I}_{v}$. These correspond to the polymatroids $(0,\log v)$ and $(\log v,0)$, respectively. While for a composite $v$, the uniform decomposition of a $\mathrm{VOA}(U_{2,3},v)$ will be more complicated. In Example 4, the uniform decomposition corresponds to the entropy function $(0.5,1.5)$ on the face $F$.
@@
$a=\frac{1}{2}H(\dfrac{|A_{i}|^{2}}{v^{2}}:i\in\mathbb{I}_{t}).$ (III.3.53)
@@
Remark Theorem III-C.2 establishes a correspondence between the face $(\hat{U}_{2,5}^{1},U^{234}_{2,3})$ characterization and suborder decomposition problem. It is obvious that $\mathrm{VOA}(U_{2,3},v)$ is inherently a suborder $\mathrm{VOA}$ of itself. On the other hand, any $\mathrm{VOA}(U_{2,3},v)$ can be decomposed into $v^{2}$ suborder VOA $\mathrm{VOA}(U_{2,3},1)$. These two cases correspond to the polymatroids $(0,\log v)$ and $(\log v,0)$, respectively. However, listing all the $\mathrm{VOA}$ decompositions of a $\mathrm{VOA}(U_{2,3},v)$ can be challenging.
@@
  • for each $i=2,3,4$, each row in $\mathbf{T}(1,i)$ occurs exactly once,
@@
$p(x_{1234})$ $=p(x_{1},x_{2},x_{3})$ (III.3.54)
$=p(x_{1},x_{2},x_{4})$ (III.3.55)
$=p(x_{1},x_{3},x_{4})$ (III.3.56)
$=p(x_{2},x_{3},x_{4})$ (III.3.57)
$p(x_{2},x_{3})$ $=p(x_{2})p(x_{3}),$ (III.3.58)
$p(x_{2},x_{4})$ $=p(x_{2})p(x_{4}),$ (III.3.59)
$p(x_{3},x_{4})$ $=p(x_{3})p(x_{4}),$ (III.3.60)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{2},x_{4})$ $=p(x_{1},x_{2})p(x_{1234}),$ (III.3.61)
$p(x_{1},x_{2},x_{3})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{3})p(x_{1234}),$ (III.3.62)
$p(x_{1},x_{2},x_{4})p(x_{1},x_{3},x_{4})$ $=p(x_{1},x_{4})p(x_{1234}).$ (III.3.63)
$p(x_{1},x_{2},x_{3})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{3})p(x_{1234}),$ (III.3.64)
$p(x_{1},x_{2},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{2},x_{4})p(x_{1234}),$ (III.3.65)
$p(x_{1},x_{3},x_{4})p(x_{2},x_{3},x_{4})$ $=p(x_{3},x_{4})p(x_{1234}).$ (III.3.66)
By (III.3.54), canceling $p(x_{1},x_{2},x_{3})$ and $p(x_{1},x_{2},x_{3},x_{4})$ on either side of (III.3.61), we have
$p(x_{1},x_{2},x_{4})=p(x_{1},x_{2}).$ (III.3.67)
Together with (III.3.55), we obtain
$p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2}).$ (III.3.68)
@@
By (III.3.57), $p(x_{1},x_{2},x_{3},x_{4})=p(x_{2},x_{3},x_{4})$, which implies that $(X_{i},i\in N_{4})$ must be uniformly distributed on the rows of a $v^{2}\times 4$ array $\mathbf{T}$ with $\mathbf{T}(2,3,4)$ a $\mathrm{VOA}(U_{2,3},v)$. Note that $p(x_{1},x_{2},x_{3},x_{4})=p(x_{1},x_{2})$ by (III.3.68), each row of $\mathbf{T}(1,2)$ occurs exactly once in $\mathbf{T}(1,2)$. Similarly, by (III.3.69), each row of $\mathbf{T}(A)$ occurs exactly once in $\mathbf{T}(A)$ for $A=\{1,3\}$ and $M=\{1,4\}$. Hence, $\mathbf{T}$ is a $\{1\}$-partial $\mathrm{VOA}(U_{2,4})$. Recall that $(X_{i},i\in N_{4})$ is uniformly distributed on the rows of $\mathbf{T}$, the probability of each row $(x_{1},x_{2},x_{3},x_{4})$ of $\mathbf{T}$ is
$p(x_{1},x_{2},x_{3},x_{4})=\frac{1}{v^{2}}.$ (III.3.70)